OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

y=mx+n m=0.4m x=5, y=-2 -2=0.4*5+n -2=2+n n=-4 So: y=0.4x-4

OpenStudy (anonymous):

x = 5 , y = -3 m = 2/5 y = mx + b -2 = 2/5 * (5) + z -2 = 2 + n n = -4 !

OpenStudy (anonymous):

thank you